E-Eyes.com Bank just issued some new preferred stock. The issue will pay a $10 annual dividend in perpetuity, beginning 4 years from now. If the market requires a 12 percent return on this investment, how much does a share of preferred stock cost today?

Far Side Corporation is expected to pay the following dividends over the next four years: $14, $11, $8, and $5. Afterward, the company pledges to maintain a constant 5 percent growth rate in dividends forever. If the required return on the stock is 13 percent, what is the current share price? (Do not round your intermediate calculations.)
Answer #1

1.Value as on year 4=Annual dividend/required return

=10/0.12

=83.33(Approx)

Hence current value=Value as on year 4*Present value of discounting factor(rate%,time period)

=83.33/1.12^3

=$59.32(Approx)

2.Value after year 4=(D4*Growth rate)/(Required return-Growth rate)

=(5*1.05)/(0.13-0.05)

=65.625

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=14/1.13+11/1.13^2+8/1.13^3+5/1.13^4+65.625/1.13^4

=$69.86(Approx).